OpenStudy (thedj4jc):

2x + 3y = 6 5x + 2y = 4 Which of the following equations could be the result of multiplication and addition to eliminate a variable in the system of equations? 19x = 24 19y = 22 11y = 22

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2x+3y=6 }\) \(\large\color{black}{ \displaystyle 5x + 2y = 4 }\) \(\large\color{black}{ \displaystyle \color{red}{(}2x+3y=6 \color{red}{)\times5} }\) \(\large\color{black}{ \displaystyle \color{red}{(}5x + 2y = 4\color{red}{)\times(-2)} }\) \(\large\color{black}{ \displaystyle 10x+15y=30 }\) \(\large\color{black}{ \displaystyle -10x - 4y = -8 }\) now add the equations.

OpenStudy (thedj4jc):

Add the equations together or work each out?

OpenStudy (solomonzelman):

add the equations together

OpenStudy (solomonzelman):

(the X is eliminated after the addition)

OpenStudy (thedj4jc):

11y = 22

OpenStudy (solomonzelman):

(( yes, so your answer is ? ))

OpenStudy (solomonzelman):

lol

OpenStudy (thedj4jc):

y = 2 Thank you!!!